1. I2+ NaClO3 ->?? 2.NH3+NaOCl->??? PLEASE ANSWER ASAP!!!This was a board question! :)

Grade:12

1 Answers

L V N Chandra Sekhar Banda
8 years ago
2NaClO3 + I2 → 2NaIO3 + Cl2

Sodium chlorate react with iodine to produce sodium iodate and chlorine. This reaction takes place in a heat, diluted solution nitric acid

Sodium hypochlorite (bleach) and ammonia do NOT yield chlorine! The reaction between the two is as followed: It does not stop with the first step.
NH3+NaOCl=NaOH+NH2Cl

NH2Cl+NaOCl=NaOH+NHCl2

NH2Cl+NaOCl=NaOH+NCl3

NH3+NH2Cl+NaOH=N2H4+NaCl+H2O